OpenStudy (anonymous):

Find all solutions to the equation. 7(sin^2)x - 14sinx+2=-5

hero (hero):

Let sinx = x: 7x^2 - 14x + 2 = -5

OpenStudy (anonymous):

so is it x=0, 2pi?

hero (hero):

Show your work.

OpenStudy (anonymous):

I honestly have no idea what I'm doing.

hero (hero):

7x^2 - 14x + 7 = 0 7(x^2 - 2x + 1) = 0 x^2 - 2x + 1 = 0 x^2 - x - x + 1 = 0 x(x - 1)-1(x - 1) = 0 (x - 1)^2 = 0 x - 1 = 0 x = 1 Replace x = sin(x) sin x = 1 x = sin^{-1}(1) x = 90

hero (hero):

x = pi/2

hero (hero):

There's only one answer for this one.

OpenStudy (anonymous):

Thank you!!